OpenStudy (anonymous):

What is the area of a square with sides of length 2x + 7? Write your answer in polynomial form.

Parth (parthkohli):

Remember - the area of a square is one side whole squared.

Parth (parthkohli):

\[ Area = (2 x + 7)^2\]

OpenStudy (anonymous):

so im not really sure what am i suppose to do?

Parth (parthkohli):

\[ (2x + 7)^2 \implies (2x + 7)(2x + 7)\]

OpenStudy (anonymous):

ooh, am i suppose to multiply?

Parth (parthkohli):

Ya

OpenStudy (anonymous):

so is that suppose 2 become 4x+49?

Parth (parthkohli):

Nope.

OpenStudy (anonymous):

hmm so what do i do?

Parth (parthkohli):

\[(2x + 7)(2x + 7) \implies 2x(2x + 7) + 7(2x + 7) \]Use the distributive property now.

OpenStudy (anonymous):

so you get 4x+14+14x+49 and then dont i do 4x+14x=18x , and then 49+14=63 so get 18x+63?

Parth (parthkohli):

4x^2*

OpenStudy (anonymous):

ooh, so is it really 18x^2+63?

Parth (parthkohli):

Nope.\[4x^2 + 14x + 14x + 49 \]What you can do is add 14x + 14x.

OpenStudy (anonymous):

to get 28x

Parth (parthkohli):

Yes.\[ 4x^2 + 28x + 49\]

Parth (parthkohli):

No wait that was correct.

Parth (parthkohli):

Yeah. It's 4x^2 + 28x + 49

OpenStudy (anonymous):

so the right answer is 4x^2+28x+49?

Parth (parthkohli):

Yep.

OpenStudy (anonymous):

okay is that all we do to it?

Parth (parthkohli):

Yes - that's it.

OpenStudy (anonymous):

okay thanks alot again :)

Parth (parthkohli):

You're welcome!

OpenStudy (theviper):

\[\Large{\color{green}{{Area \space \space of \space a \space SQUARE = side \times side}}}\]\[\Large{area=(2x+7)(2x+7)=(2x+7)^2}\]Now use the formula\[\Large{(a+b)^2=a^2+2ab+b^2\Huge{\color{red}{\checkmark}}}\]